A car leaves a certain town at noon travelling at 40 kph. At 1 pm another car leaves the town travelling at 50 kph. In how many hours will the second car overtake the first car?
--HELP!!

Answer by mananth(16946)

Car I speed 40 km/h
Car II speed 50 km/h
Since they are in same direction effective catch up speed = 50-40 = 10 kph
In 1 hour car I would have traveled 40 km
catch up time = catch up distance / catch up speed
xatch up time = 40/10= 4 hours